OpenStudy (anonymous):

A 0.145 kg baseball is thrown horizontally at 15 m/s from a platform, so that it starts at 5.2 m above the ground. How much has its kinetic energy increased when it reaches the ground? 16 J , 24 J , 0 J , or 7.4 J ? :/

OpenStudy (michele_laino):

the initial kinetic energy is: \[KE = \frac{1}{2}m{v^2} = \frac{1}{2} \times 0.145 \times {15^2} = ...Joules\]

OpenStudy (anonymous):

so we get 16.3125? so our solution is 16 J?

OpenStudy (michele_laino):

that result is right! Nevertheless it is not our final answer!

OpenStudy (anonymous):

ohh okay! what is netx?

OpenStudy (michele_laino):

we have to consider that situation:  our ball, in free falling, has gained a vertical velocity